OpenStudy (anonymous):

. Solve -4(x + 2) ≤ -9x + 6. Write your solution using set builder notation. Please show all of your work

OpenStudy (mertsj):

\[-4(x+2)\le-9x+6\] \[-4x-8\le-9x+6\] Add 9x to both sides \[5x-8\le6\] Add 8 to both sides \[5x \le14\] Divide both sides by 5 \[x \le\frac{14}{5}\]